Find all solutions of the system
y+x^2=4x
y+4x=16

Answer by mananth(16949)

y+x^2=4x
y+4x=16
y=4x-x^2
substitute y in second equation
y+4x=16
4x-x^2+4x=16
x^2-8x+16=0
(x-4)^2=0
x=4